OpenStudy (anonymous):

Given a binomial distribution with n=125 and p= 0.04, would the normal distribution provide a reasonable approximation? Why or why not? @ap statistics

OpenStudy (anonymous):

If you can walk me through it, I want to learn how to do these easier.

OpenStudy (anonymous):

The rule of thumb is usually n*p>=5 and n*(1-p)>=5. Looks like you are barely making it in this case

OpenStudy (anonymous):

n*p>=5, 125*0.04=5 n*(1-p)>=5, 125*(1-0.04)=125*0.96=120

OpenStudy (anonymous):

I don't understand, what is it close to?

OpenStudy (anonymous):

These are the criteria we use to see if a normal approximation would be good for binomial distribution. If it's true we it's reasonable

OpenStudy (anonymous):

I am sorry, but how is it true?

OpenStudy (anonymous):

So the answer to why or why not is?

OpenStudy (anonymous):

I just know the conditions. If np>=5 and n(1-p)>=5, a normal approximation is reasonable.